OpenStudy (anonymous):

Solve -8-11(3x +7≤-40x-9 write solution using set builder notation please show work!

OpenStudy (anonymous):

Its (3x+7)

OpenStudy (anonymous):

Why is the parenthesis not closed ?

OpenStudy (anonymous):

Oh, never mind.

OpenStudy (anonymous):

i will reposted it!

OpenStudy (anonymous):

I got: \[\left\{ x \in \mathbb{R} | x \le \frac{76}{7} \right\} \]

OpenStudy (anonymous):

You just solve for x. Distribute:\[-8-33x-77 \le -40x-9\] \[-33x-85 \le -40x-9\] \[7x \le 76\] \[x \le \frac{76}{7}\] The beginning of the set just tells you that its any real number (which it is)

OpenStudy (anonymous):

Alternative form : \[\left \{ x \in \mathbb{R} | x \in ]- \infty ; 76/7 ] \right\}\]